### v15

# Insuring uninsurable income

Abstract
We study dynamic mechanism design in a pure-exchange economy with privately observed idiosyncratic income. Classic hidden-income contracts attain constrained efficiency only at the cost of immiseration (Green 1987; Thomas–Worrall 1990). We propose a simple recursive mechanism—adapted from Marcet–Marimon (1992)—that shifts each income shock forward by one period, keeps promised utilities in a bounded set, and, under a transparent “moderate risk-aversion” condition, delivers sequential efficiency. In a stationary overlapping-generations setting, we further provide an explicit condition on the initial promise that ensures budget sustainability; early cohorts pre-fund intertemporal smoothing so that every cohort attains a higher expected lifetime utility than under autarky. Our analysis uses a single state (promised utility), closed-form transfers, and a Bellman verification.
Keywords: Dynamic mechanism design; recursive contracts; private information; immiseration; overlapping generations.
JEL: D82, D86.
1 Introduction
Private information can render idiosyncratic income effectively uninsurable. Since the seminal work of Green [1987] and Thomas and Worrall [1990], the constrained-efficient hidden-income contract trades insurance off against incentives and features a drift of average promised utilities toward a lower bound—immiseration—with cross-sectional dispersion rising over time. [^2] Earlier papers sometimes use immiserization[e.g., Phelan, 1998, Zhang, 2009]. Recent work with persistent private information, Bloedel et al. [2025], also sharpens the mechanism of immiseration and explains the balance of the cost of incentive provision and backloading of high-powered incentives. When agents are almost perfectly patient, the drift can vanish [Carrasco et al., 2019], but the knife-edge offers little comfort for moderate discounting.
A different route was proposed by Marcet and Marimon [1992]. Allowing agents to make productivity-enhancing investments, they designed a transfer rule that shifts risk across periods for the same individual. Capital accumulation enlarges the future consumption set and breaks the immiseration trap; the associated recursive saddle-point representation, later formalized by Marcet and Marimon [2019] makes the analysis more tractable. Yet the mechanism has not been explored in a pure-exchange economy without investment or storage, where the scope for avoiding downward drift is a priori unclear.
This paper adapts the Marcet–Marimon risk-shifting mechanism to an exchange economy in which agents cannot save. Promised utility is the sole state variable, so the model yields a closed-form characterization. If individual assets were allowed, one would require the set-valued approximation methods of Sleet and Yeltekin [2016]—a contrast confirmed by our closed-form example in Section 4.
The contribution of my paper is threefold. (i) A recursive, one-state dynamic mechanism that postpones shocks one period ahead and eliminates immiseration while preserving sequential incentive compatibility; (ii) a transparent sufficient condition for sequential efficiency under moderate risk aversion; (iii) in a stationary overlapping-generations society, an explicit inequality on the initial promise that suffices for budget sustainability, with a simple economic interpretation; early cohort pre-fund later transfers; the mechanism therefore delivers higher expected lifetime utility than autarky for all cohorts.
The organization of the paper is as follows. Section 2 sets up the model and transfer rule. Section 3 proves efficiency and boundedness, and analyzes inter-generational sustainability. Section 4 presents numerical illustrations. Section 5 verifies optimality via a Bellman equation, and Section 6 concludes.
2 Model
2.1 Environment
We consider a risk-neutral planner allocating transfers across a continuum of risk-averse individuals indexed by birth cohorts. Individuals survive to the next period with probability $\alpha\in(0,1)$ and discount at rate $r\in(0,1)$; their effective discount factor is $\beta=\alpha r$. In each period, a newborn cohort enters so that the cross-section is stationary. [^3] A similar stationary population model is used in Fujiwara-Greve and Okuno-Fujiwara [2009] for their Prisoner’s Dilemma game. Each individual draws an i.i.d. income shock $e_{t}\in E:=\{e^{1},e^{2},\dots,e^{M}\}\subset\mathbb{R}_{+}$ with $e^{i}<e^{j}$ for $i<j\in\{1,\dots,M\}$, $M\geq 2$. The planner commits to a recursive transfer mechanism $\mathcal{M}$ mapping the current promised $\lambda_{t}>0$ into a transfer $\tau_{t}$ and next period’s promise $\lambda_{t+1}$. The single state variable is the promised utility $\overline{v}_{1}(\lambda)$, which is strictly increasing in $\lambda$.
2.2 Symmetric-information benchmark
First, we derive an optimal transfer function which, under symmetric income information, ensures a level of utility corresponding to a promise $\lambda_{0}>0$ . Following Marcet and Marimon [1992], we will use this promise $\lambda_{0}$ as a state variable under asymmetric income information. For this purpose, in this subsection only, we consider the case where individuals cannot hide their income from the planner in each period.
At period 0, the planner is considering an efficient transfer mechanism that increases individuals’ lifetime utility as much as possible, but it cannot afford an infinite net subsidy. The planner solves the following problem, which is normalized by multiplying $(1-\beta)$, having decided on an upper bound for the net subsidy.
$(\ast)\quad\max_{\{\tau_{t}\}_{t=0}^{\infty}}(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})\right]\quad\text{s.t. }\quad(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}(\tau_{t})\right]\leq\overline{C},$
where $\overline{C}>0$ is an exogenous upper bound on the present value of net subsidies, $\mathbb{E}_{0}$ is the conditional expectation given the information in period 0, $u:\mathbb{R}_{+}\to\mathbb{R}$ is the individual’s utility function, $c_{t}=e_{t}+\tau_{t}$ is the consumption in period $t$ starting from the individual’s birth period. The utility function $u$ is assumed to be $u^{\prime}>0$, $u^{\prime\prime}<0$, to satisfy the Inada conditions ($\lim_{c\to 0}u^{\prime}(c)=+\infty$, $\lim_{c\to\infty}u^{\prime}(c)=0$) and to be bounded from above.
Without specifying $\overline{C}$, $\{\tau_{t}\}_{t=0}^{\infty}$ is determined by $\lambda_{0}$, and the same $\lambda_{0}$ determines $\overline{C}$ as well. Let $\mu>0$ be the multiplier on the subsidy constraint and put $\lambda_{0}=1/\mu$. The problem $(\ast)$ can be rewritten as the following problem.
$\max_{\{\tau_{t}\}_{t=0}^{\infty}}$ $(1-\beta)\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\beta^{t}[\lambda_{0}u(c_{t})-\tau_{t}]\right].$ (1)
From the first order condition,
$u^{\prime}(c_{t})=\frac{1}{\lambda_{0}}\text{ for }t\in\mathbb{N}\cup\{0\}$
and the optimal policy delivers a constant consumption $c^{\ast}(\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})$ and transfer $\tau^{\ast}(e|\lambda_{0})=(u^{\prime})^{-1}(\lambda_{0}^{-1})-e$, where $\tau^{*}(\cdot\ |\lambda_{0})$ is a time-invariant transfer function. We record two objectives for future use: Let $v_{1}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing each individual’s lifetime utility provided by the transfer mechanism. Let $v_{2}:\mathbb{R}_{++}\times\mathbb{R}_{+}\to\mathbb{R}$ be a function expressing the expected balance of transfers. Then, for given $\lambda_{0}>0$ and $e_{0}\in E$, these values are written as follows.
$v_{1}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}u(c_{t})]=u((u^{\prime})^{-1}(\lambda_{0}^{-1}))=:\overline{v}_{1}(\lambda_{0})\,\text{(promised utility)}$
$v_{2}(\lambda_{0},e_{0})$ $:=(1-\beta)\mathbb{E}_{0}[\sum_{t=0}^{\infty}\beta^{t}(-\tau_{t})]\,\text{(expected budget impact)}.$
The map $\lambda\mapsto\overline{v}_{1}(\lambda)$ is strictly increasing and will serve as the state variable once income becomes private. We also write
$\overline{v}_{2}(\lambda_{t}):=\mathbb{E}[(1-\beta)\mathbb{E}_{t}[\sum_{s=t}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))]].$
Remark 1.
(Lifetime value functional) Let $\mathcal{T}$ be the set of all admissible control functions. Given an initial promise $\lambda_{t}$ and any admissible transfer rule $\tau\in\mathcal{T}$, each individual’s discounted lifetime utility is
$J(\lambda_{t},t,\tau)=(1-\beta)\mathbb{E}_{t}\left[\sum_{k=t}^{\infty}\beta^{k-t}u(e_{k}+\tau(e_{k}|\lambda_{t}))\right]\quad\text{for }\tau\in\mathcal{T}.$ (2)
This function will be used in Section 5 to verify that our mechanism solves the associated Bellman equation.
2.3 Private information and mechanism design
If income is private information, the planner has to ask individuals to declare their income. In general, a mechanism $\Gamma$ that arranges transfers $\{\tau_{t}\}_{t=0}^{\infty}$ needs to satisfy sequential incentive compatibility defined below.
Definition 1.
[Sequential incentive compatibility, Marcet and Marimon, 1992] Given the current promise $\lambda$, truthful reporting maximizes the continuation value. Equivalently,
$(1-\beta)u(e+\tau(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\geq(1-\beta)u(e+\tau(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\quad\forall e,\tilde{e}\in E.$
We further evaluate a mechanism according to the following notion.
Definition 2.
[Sequentially efficient mechanism, Marcet and Marimon, 1992] The mechanism $\Gamma$ is said to be a sequentially efficient mechanism if it is sequentially incentive compatible and it is not Pareto dominated by any other sequentially incentive compatible mechanism.
In addition to taking these things into account, since it is known to lead to an immiseration outcome, the planner does not hope to manipulate the transfer function $\tau^{*}$. The planner takes an alternative approach. It is to vary the promise $\lambda$ in each period. This could be an incentive to tell the truth. For a given $\lambda$ in a given period, the planner promises to ensure that the lifetime utility level $\overline{v}_{1}(\lambda)$, and the promise for the next period is renewed taking into account the current promise and the income reported in the current period. If an individual truthfully reports income $e$ in a period $t$, then the lifetime utility from that period should be
$\overline{v}_{1}(\lambda)=\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))\big{]},$ (3)
where $\lambda^{\prime}(\cdot|\lambda):E\to\mathbb{R}_{++}$ is the promise for the next period. If $\lambda^{\prime}$ satisfies (3), then $\mathbb{E}[\overline{v}_{1}(\lambda^{\prime}(e|\lambda)]=\overline{v}_{1}(\lambda)$. Truth-telling is then induced if, given the current promise $\lambda$, the renewed promise $\lambda^{\prime}$ satisfies the following incentive constraints.
$(1-\beta)u(e+\tau^{*}(e|\lambda))$ $+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
$\geq(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (4)
Such a promise $\lambda^{\prime}$ that satisfies both (3) and (4) could be adapted from the $\lambda$ mechanism of Marcet and Marimon [1992], provided that there is a feasible $\lambda^{\prime}$. By a feasible $\lambda^{\prime}$ we mean that the updated promise keeps the level of expected utility for the next period below the supremum of each individual’s utility function. That is, let $S$ be the supremum of each individual’s utility function, $S=\sup_{x\in\mathbb{R}_{+}}u(x)$. For given $\lambda>0$, if $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))<S$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\big{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e)-\overline{v}_{2}(\lambda))\big{)}.$ (5)
Under the updating rule of $\lambda^{\prime}$ in (5), misreporting never raises the weighted sum $\lambda u(e+\tau^{*}(e|\lambda))+v_{2}(\lambda,e)$, hence truth-telling is optimal.
On the other hand, if there is an income $\hat{e}\in E$ for given $\lambda$ such that $\lambda^{\prime}$ defined in (5) must promise the expected utility for the next period above the supremum of the individual’s utility function, then the planner must find the closest promise for which there is a certainty equivalent to the promised level of lifetime utility. That is, for given $\lambda$, if there is $\hat{e}\in E$ such that $\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\hat{e})-\overline{v}_{2}(\lambda))\geq S$, for reported income $e\in E$, define $\lambda^{\prime}$ such that
$\lambda^{\prime}(e|\lambda)=\overline{v}_{1}^{-1}\bigg{(}\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})-\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]\bigg{)},$ (6)
where $\mathbb{E}_{\tau^{\ast}}$ denotes
$\mathbb{E}_{\tau^{\ast}}\left[v_{2}(\lambda,\min\{e,\overline{e}_{\lambda}\})\right]:=(1-\beta)\mathbb{E}\left[-\tau^{\ast}(\min\{e_{t},\overline{e}_{\lambda}\}|\lambda)+\sum_{s=t+1}^{\infty}\beta^{s-t}(-\tau^{\ast}(e_{s}|\lambda_{s}))\right]$
$\overline{e}_{\lambda}$ is defined such that
$\overline{e}_{\lambda}=\arg\max_{e^{\prime}\in E}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}$
subject to
$\overline{v}_{1}(\lambda)+\frac{1}{\lambda\beta}\big{(}v_{2}(\lambda,e^{\prime})-\mathbb{E}_{\tau^{\ast}}[v_{2}(\lambda,\min\{e_{t},e^{\prime}\})]\big{)}<S.$
This adjustment also changes the transfers in the current period $t$. However, we see the recursive relation in (6) satisfies the following corresponding incentive constraint.
$(1-\beta)u(e$ $+\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(e|\lambda))$
$\geq(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda}\}|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime}(\tilde{e}|\lambda))\text{ for }e,\tilde{e}\in E.$ (7)
This is because the transfer $\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda)$ in (7) is the solution to the problem (1) subject to $\tau(e_{t})\geq(u^{\prime})^{-1}(\lambda^{-1})-\overline{e}_{\lambda}$ for $t\in\mathbb{N}\cup\{0\}$, where $\lambda=\lambda_{0}$. In other words, since the transfer to the planner is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
Concerning the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (7) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
$\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
3 Risk-shifting mechanism
This section builds and analyzes the transfer rule $\mathcal{M}$. We first show that each individual’s promised utility forms a bounded martingale, and then derive conditions under which the planner’s inter-temporal budget is non-negative.
Since the definition of $\lambda^{\prime}$ in (6) is an extension of (5), we can use it for the case where $\lambda^{\prime}$ could also be defined in (5). We define a mechanism like this.
Definition 3.
(Risk-shifting mechanism) A mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6). Within each period $t$, consumption is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1})$; schocks affect only $\lambda_{t+1}$.
A sufficient condition for the sequence of promise $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (5) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal consumption $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$.
Assumption 1.
(Moderate risk aversion). For all $\lambda>0$,
$-\frac{u^{\prime\prime}}{u^{\prime}}((u^{\prime})^{-1}(\lambda^{-1}))<\frac{\beta}{(1-\beta)(e^{M}-\mathbb{E}[e_{t}])}.$ (8)
Interpretation. Arrow-Pratt risk aversion at the contract consumption level is smaller than a scaled income spread; the condition is mild when $\beta$ is close to one. Under Assumption 1, feasibility does not bind in the promise-update map, and the recursive rule (5) applies period by period.
Remark 2.
We could remove condition Assumption 1 by removing the assumption of bounded utility; dropping it would not violate sequential efficiency but would complicate the martingale proof. Boundedness is used to ensure that $\{\overline{v}_{1}(\lambda_{t})\}$ converges.
Lemma 1.
Under Assumption 1 the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5).
Proof: see Appendix A.
If the inequality (8) does not hold, there may be a case where a promised-utility state $\lambda_{t}$ in $\mathcal{M}$ is defined in (6) instead of in (5), e.g. if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathbb{E}[e_{t}]))$.
The risk-shifting mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (5) or in (6). Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
Theorem 1.
The mechanism $\mathcal{M}$ defined by $(\tau^{\ast},\{\lambda_{t}\})$ is sequentially incentive compatible. Under Assumption 1, it is sequentially efficient.
Proof: see Appendix A.
3.1 Individual level: bounded promised utilities
This subsection shows that the sequence of promised utilities forms a bounded martingale and therefore converges, so immiseration cannot occur.
Let $(\mathcal{F}_{t})$ be the natural filtration generated by $\{\lambda_{s}\}_{s\leq t}$. Because (6) makes $\lambda_{t+1}$ measurable w.r.t. the $\sigma$-algebra generated by $\lambda_{t}$ and $e_{t}$, the process $\overline{v}_{1}(\lambda_{t})$ is adapted.
Theorem 2.
The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a bounded martingale and converges to an integrable random variable a.s. (no immiseration).
Proof: see Appendix A.
Because promised utility is a martingale with bounded support, the mechanism is a fair intertemporal lottery: expected lifetime utility is preserved ex-ante, but its realisation drifts neither up nor down on average.
3.2 Planner’s budget and sustainability
All claims in this subsection are sufficient (not necessary) conditions based on a second-order expansion. We derive sufficient conditions under which the cross-sectional budget is non-negative each period.
The convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism. This is because the planner predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma. Let $\overline{\tau}_{n}$ denote the average transfer to an $n$-period-old individual (positive means the planner pays). Then $\overline{\tau}_{n}$ is:
$\overline{\tau}_{n}$ $=\mathbb{E}[\tau^{\ast}(e_{n}|\lambda_{n})]=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
Lemma 2.
The average transfer for each cohort $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
Proof sketch (details in Appendix B). For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
The proof of Lemma 2 shows that the increasing trend in $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is due to the cost of keeping the expected lifetime utility in each period unchanged from the previous period, although there are risks between periods.
Let $g_{t}(\lambda_{0})$ be the expected budget balance in period $t$ for the planner for a given initial promise $\lambda_{0}$. Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of continuum of individuals on the $(1-\alpha)$ interval is born in each period, $g_{t}(\lambda_{0})$ is given by
$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
Here, a positive $g_{t}$ denotes a surplus (the planner collects funds). For the $\mathcal{M}$ mechanism to be sustainable, the budget balance in each period must be greater than or equal to zero. As we see in the following proposition, under a second-order (Taylor) approximation, if incomes have a symmetric distribution, a sufficient condition for the budget balance in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
Theorem 3.
In a stationary overlapping-generations society with i.i.d. income risk, if
  • (i)
    incomes have a symmetric distribution,
  • (ii)
    $u^{\prime\prime}/u^{\prime}$ is a decreasing function,
  • (iii)
    $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function and
  • (iv)
    initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied,
then planner’s period-by-period balance $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ under a second-order (Taylor) approximation.
Proof: Second-order approximation shows decreasing differences. See Appendix C.
4 Numerical examples
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Theorem 3 is given by the second-order polynomial approximation as follows.
Corollary 1.
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Theorem 3 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
$\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))\geq\sqrt{\frac{(1-\alpha r)^{2}}{\alpha r^{2}(1-\alpha)}\mathrm{Var}[e_{t}]},$ (9)
where $\lambda_{E}=1/(u^{\prime}(\mathbb{E}[e_{t}]))$ denotes the promise corresponding to mean income report.
Proof: see Appendix D.
Figure 1: Utility Function
{singlespace*}
Notes: This figure illustrates a utility function (vertical axis) against consumption (horizontal axis). The segment $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ appears in Corollary 1. It is the linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
The left-hand side of the inequality (9) is a linear approximation of the difference between the expected income and the consumption level in the mechanism $\mathcal{M}$ with the promise $\lambda_{0}$. The inequality (9) means that this length is greater than the square root of the weighted income variance. Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$. The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ is higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (9).
A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$ for making the mechanism $\mathcal{M}$ sustainable. If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (9) may be unnecessarily strict.
We illustrate the mechanism with a CRRA utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
  • Utility $u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma}$, $\gamma=-3$
  • Income support $E=\{0,1,2,3,4,5,6,7,8,9\}:=E_{s}$
  • Probabilities $P(e)$ (mirror-symmetric, approx. normal) $P_{s}$
  • Discount factor and survival rate $r=\alpha=0.93$
  • Initial state $\lambda_{0}=50\in[\lambda_{CE},133]$
The values follow Corollary 1: the lower bound $\lambda_{CE}=1/u^{\prime}\bigl{(}E[u(e_{t})]\bigr{)}=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals of the same generation and corresponding transfer balances for $100$ periods for the above case. The initial promise is set $\lambda_{0}=50$. Consistent with Theorem 2, the utility path stays bounded while transfers $\{-\overline{\tau}_{n}\}_{n=0}^{100}$ fall, echoing Lemma 2.
Figure 2: Sample mean of lifetime utilities and transfer balances
Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. In this society, each individual will be alive in the next period with probability $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$. As in Figure 2, each individual’s utility is $u(c)=(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial promise $\lambda_{0}=50$ and the discount factor $r=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated with the same symmetric distribution $P_{s}$.
The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and supports Theorem 3.
Figure 3: Sample means of budget balance $\{g_{t}(\lambda_{0})\}$
Remark 3.
Classic hidden-income contracts [e.g. Green, 1987, Thomas and Worrall, 1990] spread shocks within each period: truth telling is rewarded by a higher spot transfer, so lifetime utilities diverge over time and immiseration follows. Under our mechanism, by contrast, the realized income $e_{t}$ only determines next period’s promise $\lambda_{t+1}$; consumption in period $t$ is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda^{-1}_{t})$ with probability $1$. The shock is therefore shifted one step forward—a transfer of risk to the future. This intertemporal insurance does make the average transfer profile $\{-\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ fall with age, threatening the planner’s long-run budget. Theorem 3 shows the problem can be neutralized by choosing an initial promise $\lambda_{0}$ that gives early cohorts negative net transfers, so later positive transfers are prefunded. Finally, note a key difference from Marcet and Marimon [1992]: in their growth model, investment returns are themselves the incentive, and the contract is bilateral and non-competitive, which can depress the manager’s utility below autarky. Our pure-exchange setting has no such external scale effect, so every cohort attains a higher lifetime utility than autarky.
5 Dynamic-programming verification
This section verifies that the transfer rule $\tau^{\ast}$ solves the infinite-horizon Bellman equation and that $V$ is the corresponding value function (under condition (8) of Lemma 1).
The Bellman equation (DP) is given by
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)]\},$ (10)
where $\Lambda(\lambda_{t})$ is a set of measurable controls satisfying incentive and promise-keeping, and $\lambda_{t+1}$ is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{R}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (6).
Our candidate solution is $V$ in (11) with control $\tau^{\ast}$.
$V(\lambda_{t},t)=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]},\quad\text{for }\,\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}.$ (11)
We observe the following proposition.
Proposition 1.
With Assumption 1, $V$ in (11) solves the infinite-horizon Bellman equation, and $\tau^{*}$ maximizes the lifetime value $J$ in (2).
Proof sketch.
[…]
(complete proof in Appendix E).
6 Conclusion
This paper shows that the grim ”immiseration“ prediction of Green [1987] and Thomas and Worrall [1990] is not inevitable. A simple risk shifting rule that postpones risk can deliver near full within-period insurance while keeping promised utilities bounded. Provided the initial promise satisfies a transparent inequality, inter-generational cooperation sustains the mechanism, and every cohort attains higher lifetime utility than under autarky, with no within-cohort inequality.
Appendix A Proof of propositions 1 and 2
Proof of Lemma 1.
Write $F(\lambda)=\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))$ for $\lambda>0$. It suffices to show the proposition that $F(\lambda)<S$ for all $\lambda>0$.
Since the utility function $u$ is continuous, bounded from above and satisfies the Inada condition, we have
$\lim_{\lambda\to\infty}\overline{v}_{1}(\lambda)=u(\lim_{\lambda\to\infty}(u^{\prime})^{-1}(\lambda^{-1}))\leq S$
$\lim_{\lambda\to\infty}(u^{\prime})^{-1}(\lambda^{-1})=\infty.$
So we see that
$\lim_{\lambda\to\infty}F(\lambda)=\lim_{\lambda\to\infty}\overline{v}_{1}(\lambda)=S.$
If (8) holds, then we have $F^{\prime}(\lambda)>0$ for $\lambda>0$, which implies that $F$ is strictly increasing in $\lambda$. Consequently, we see the desired result $F(\lambda)<S$ for all $\lambda>0$. ∎
Proof of Theorem 1.
With respect to $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ defined throughout in (5), the propositional statement is shown by Marcet and Marimon [1992]. For completeness, we report the entire proof.
We first show that $\mathcal{M}$ is sequentially incentive compatible. For $\lambda_{t}>0$ with $F(\lambda_{t})<S$
$\lambda_{t}\big{[}$ $(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e|\lambda_{t}))\big{]}$
[…]
$\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},\tilde{e})-\overline{v}_{2}(\lambda_{t})$
$=\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\tilde{e}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(\tilde{e}|\lambda_{t}))\big{]}\text{ for }e,\tilde{e}\in E.$
The last inequality follows from the optimality of $\lambda\overline{v}_{1}(\lambda)+v_{2}(\lambda,e)$ in the problem (1) given $\lambda$. Hence, $\mathcal{M}$ is sequentially incentive compatible for $\lambda>0$ defined in (5).
For $\lambda_{t}>0$ with $F(\lambda_{t})\geq S$
$\lambda_{t}\big{[}(1-\beta)$ $u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(e|\lambda_{t}))\big{]}$
[…]
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
$\geq\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
$=\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{\tilde{e},\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t}^{\prime}(\tilde{e}|\lambda_{t}))\big{]}\text{ for }e,\tilde{e}\in E.$
The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) (planner’s static problem at given $\lambda$) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
We proceed to show that if condition (8) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (5). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
It remains to prove that under condition (8) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$. Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$. Since condition (8) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1. We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (8). Then, by construction, each individual has the same present value for both contracts. Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$. However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under the condition (8). ∎
Proof of Theorem 2.
We first show that $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale. Since the sequence $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}\in\mathcal{M}$ satisfies (6), we have
$\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
for given realized $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{B}\overline{v}_{1}(\lambda_{n-1})(\omega)dP_{\overline{v}_{1}(\lambda_{n-1})},$
and
$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n})dP$
$=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n-1})dP.$
From the last equality we have for $n\in\mathbb{N}$ $\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e. $[P]$, and $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale as claimed.
We proceed to show that $\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$. From (5) and (6) we have $\overline{v}_{1}(\lambda_{n})\leq S$ for $n\in\mathbb{N}$. So we have $\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$.
Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathbb{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere. ∎
Appendix B Proof of Lemma 2
Proof of Lemma 2.
We first consider the case $n=1$. For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
For $n>1$, since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
$\overline{\tau}_{n-1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathbb{E}[e_{n-1}]$
$=\mathbb{E}[u^{-1}(\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathbb{E}[e_{n-1}]$
$\leq\mathbb{E}[\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathbb{E}[e_{n-1}]$
$=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]$
$=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
Appendix C Proof of Theorem 3
The proof of Theorem 3 uses the following lemma.
Lemma C.1.
Under the assumptions (i), (ii), (iii) in Theorem 3, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.
Proof.
First we consider the case where $\{\lambda_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ can be defined in (5). Using the third-order Taylor polynomial, for a given promise $\lambda_{n-1}>0$ and a given income $e\in E$, the corresponding transfer is written as
$u^{-1}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))-u^{-1}(\overline{v}_{1}(\lambda_{n-1}))=$ $\frac{1}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\Delta_{n-1}-\frac{1}{2!}\frac{u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{3}}\Delta_{n-1}^{2}$
$+\frac{1}{3!}\left(-\frac{u^{\prime\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{4}}+\frac{3u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{2}}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))^{5}}\right)\Delta_{n-1}^{3}$
$+h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))\Delta_{n-1}^{3},$ (C.1)
where $\Delta_{n-1}=(1-\beta)(e-\mathbb{E}[e_{n}])/(\beta\lambda_{n-1})$ and $h_{3}:\mathbb{R}\to\mathbb{R}$ is a function such that $\lim_{\Delta_{n-1}\to 0}h_{3}(\overline{v}_{1}(\lambda_{n}(e|\lambda_{n-1})))=0$. Taking the expectation of (C.1), $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is written as
$\overline{\tau}_{n}-\overline{\tau}_{n-1}$ $=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}[e_{n}]\right]+o(\Delta_{n-1}^{3}).$
Note that since incomes follow a symmetric distribution, the third term on the right-hand side of the equation (C.1) will cancel out when the expectation is taken.
Since $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function, from Jensen’s inequality we have
$\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\left(u^{-1}(\overline{v}_{1}(\lambda_{n-1}))\right)\right]$ $=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}\big{(}\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})]\big{)}\big{)}\right]$
$\geq\mathbb{E}\left[\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n}))\big{)}\bigg{|}\overline{v}_{1}(\lambda_{n-1})\right]\right]$
$=\mathbb{E}\left[-\frac{u^{\prime\prime}}{u^{\prime}}\big{(}u^{-1}(\overline{v}_{1}(\lambda_{n})))\right].$
It follows that the third order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence.
We can now proceed analogously to the proof of the case where $\lambda_{n-1}$ must be defined in (6). We change $\Delta_{n-1}$ in (C.1) to
$\overline{\Delta}_{n-1}=\frac{(1-\beta)\big{(}\min\{e,\overline{e}_{\lambda_{n-1}}\}-\mathbb{E}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\big{)}}{\beta\lambda_{n-1}}.$
By the asymmetric distribution of conditional income $e\leq\overline{e}_{\lambda_{n-1}}$, the third term on the right-hand side of (C.1) remains and $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is given by
$\overline{\tau}_{n}-\overline{\tau}_{n-1}=\mathbb{E}\left[\frac{1}{2!}\left(\frac{-u^{\prime\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}{u^{\prime}(u^{-1}(\overline{v}_{1}(\lambda_{n-1})))}\right)\left(\frac{1-\beta}{\beta}\right)^{2}\mathrm{Var}\big{[}\min\{e,\overline{e}_{\lambda_{n-1}}\}\big{]}\right]+o(\overline{\Delta}_{n-1}^{2}).$
By a similar argument, the second-order polynomial approximation of $\{\overline{\tau}_{n}-\overline{\tau}_{n-1}\}_{n\in\mathbb{N}}$ is a decreasing sequence, and the proof is complete. ∎
We prove Theorem 3 below.
Proof of Theorem 3.
From Lemma C.1, with the second-order polynomial approximation, we have
$\overline{\tau}_{n}-\overline{\tau}_{n-1}\leq\overline{\tau}_{1}-\overline{\tau}_{0}\quad\text{for }n\in\mathbb{N}.$
Hence, the following inequality holds.
$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}$
$\geq\alpha\cdot\frac{1-\alpha^{t}}{1-\alpha}(\overline{\tau}_{0}-\overline{\tau}_{1})-\overline{\tau}_{0}.$
The last term is positive if
$\overline{\tau}_{0}\geq\frac{\alpha(1-\alpha^{t})}{1-\alpha}(\overline{\tau}_{1}-\overline{\tau}_{0})$
and we see a sufficient condition for $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ is $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
Appendix D Proof of Corollary 1
Proof of Corollary 1.
Let us first examine the case where $\lambda_{1}$ can be defined in (5). Applying Taylor theorem to $u^{-1}(\overline{v}_{1}(\lambda_{1}))$, $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)$ is expressed as
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{0}]+o(\Delta_{0}^{3}).$ (D.1)
On the other hand, applying Taylor to $u^{-1}(\overline{v}_{1}(\lambda_{E}))$, we have
$-\overline{\tau}_{0}$ $=-u^{-1}(\overline{v}_{1}(\lambda_{0}))+u^{-1}(\overline{v}_{1}(\lambda_{E}))$
$=\frac{(u^{-1})^{\prime}(\overline{v}_{1}(\lambda_{0}))}{1!}\Delta_{E}+\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\Delta_{E}^{2}+o(\Delta_{E}^{2}),$ (D.2)
where $\Delta_{E}=\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0})$. Since the first term on the right-hand side of (D.2) is positive, it is sufficient to show the second term on the right-hand side of (D.2) is equal to or greater than the first term on the right-hand side of (D.1). However, from the corollary assumption (9), we have
$\Delta_{E}^{2}\geq\frac{(1-\beta)^{2}}{\lambda_{0}^{2}\beta^{2}}\frac{\alpha}{1-\alpha}\mathrm{Var}[e_{t}],$
which is the desired conclusion that the condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ holds with the second-order polynomial approximation. For the case where $\lambda_{1}$ must be defined in (6), we have
$\frac{\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})}{1-\alpha}$ $=\frac{(u^{-1})^{\prime\prime}(\overline{v}_{1}(\lambda_{0}))}{2!}\left(\frac{1-\beta}{\beta\lambda_{0}}\right)^{2}\frac{\alpha}{1-\alpha}\mathrm{Var}\big{[}\min\{e_{0},\overline{e}_{\lambda_{0}}\}\big{]}+o(\overline{\Delta}_{0}^{2}).$
Since $\mathrm{Var}[\min\{e_{0},\overline{e}_{\lambda_{0}}\}]\leq\mathrm{Var}[e_{0}]$, (9) is sufficient to be $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$. ∎
Appendix E Proof of Proposition 1
Proof of Proposition 1.
We observe first that if (8) holds, then $V$ defined in (11) satisfies DP. Indeed, since by definition of $\tau^{*}$ we have $u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))=\overline{v}_{1}(\lambda_{t})$,
$V(\lambda_{t},t)$ $=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}\big{[}\mathbb{E}_{t+1}\big{[}\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}\big{]}$
$=(1-\beta)u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.1)
But $\tau^{*}$ is the solution to the problem
$\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)].$ (E.2)
Indeed, for given information in period $t$, the integrand of the objective function in (E.2) multiplied by $\lambda_{t}$ is written as
$\lambda_{t}\big{[}(1-\beta)$ $u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t+1}(e_{t}|\lambda_{t}))\big{]}$
$=\lambda_{t}\big{[}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e_{t})-\overline{v}_{2}(\lambda_{t})$
and we see that the last terms are maximized by $\tau^{*}$. It follows that (E.1) is written as
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)],$
and we see the desired result.
The task is now to check the terminal condition. Since $u$ is bounded from above, $\lim_{t\to\infty}V(\lambda_{t},t)\beta^{t}=0$. Hence, the sufficient terminal condition of Wiszniewska-Matyszkiel [2011:  Thm.1] is met, and we see $V$ defined in (11) is the value function of the Bellman equation in DP, and $\tau^{*}$ maximizes $J$ in (2). ∎